For the IR spectrum of ethanol, match the motion to the wavenumber.

sapling learning For the IR spectrum of ethanol, match the motion to the wavenumber. 3400 cm-1 2900 cm-1 1450 cm-1 1050 cm-1 O-H stretch H-C-H bend C-H stretch C-O stretch
Answer #1

Alcohols have characteristic IR absorptions for O-H and the C-O stretching vibrations.

O-H= 3400 cm-1

C-H Stretch = 2900 cm-1
H-C-H bend = 1450 cm-1

C-o Stretch = 1050 cm-1